Clifford, Inc., has a target debt–equity ratio of .81. Its WACC is 8.5 percent, and the tax rate is 34 percent.

a. If the company’s cost of equity is 12.1 percent, what is its pretax cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
  
Pretax cost of debt             %

b. If the aftertax cost of debt is 5.2 percent, what is the cost of equity? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
  
Cost of equity             %
Answer #1

a. WACC = Weight of debt * After tax cost of debt + Weight of equity * Cost of equity

8.5% = 0.81 / 1.81 * After tax cost of debt + 1/ 1.81 * 12.1%

After tax cost of debt = 0.040556

Pretax cost of debt = 0.040556 / ( 1- 34%) = 0.0614 or 6.14% (rounded off)

b.

WACC = Weight of debt * After tax cost of debt + Weight of equity * Cost of equity

8.5% = 0.81 / 1.81 * 5.2% + 1/ 1.81 * Cost of equity

Cost of equity = 11.17%

The figures may slightly vary due to rounding off differences.